OpenStudy (konradzuse):

find the equation of the tangent line of y = 3x2 − ln(x) for (1,3)

OpenStudy (konradzuse):

Not sure how to set this up.

terenzreignz (terenzreignz):

Calculus is needed here, are you up for it?

OpenStudy (konradzuse):

yeah I'm always ready :P

terenzreignz (terenzreignz):

Ok, first step is to find dy/dx, this is the slope of the tangent line at any given point

OpenStudy (konradzuse):

It says the answer is 5x-2. So I figured deriving would be 6x - 1/x

OpenStudy (konradzuse):

plugging in x would be 6* 1 - 1/1 = 5

terenzreignz (terenzreignz):

yeah, so 5 is the slope of the line and we have a point (1,3) so all that's left is to determine the equation of the line from the point and the slope

terenzreignz (terenzreignz):

so you get y - 3 = m(x - 1) y - 3 = 5x - 5 y = 5x - 2 hmm I'm slow haha

OpenStudy (anonymous):

Complete step-by-step slope at x=1 (given by (1,3) \[\frac{d}{dx} y = \frac{d}{dx} 3x^2 - \frac{d}{dx} \ln(x)\] applying power rule and chain rule \[\frac{d}{dx} y = \frac{d}{dx} (3)(2)x^{2-1} - \frac{d}{dx} \frac{1}{ln(e)*x}\] \[\frac{dy}{dx} = 6x - \frac{1}{x}\] \[\frac{dy}{dx} = 6(1) - \frac{1}{(1)}=5\] \[y - y_0 = m(x - x_0) \] \[y - 3 = m(x - 1) \] \[y - 3 = 5x - 5 \] \[y = 5x - 2\]

OpenStudy (konradzuse):

Sorry for not replying, my internet dropped again :(

OpenStudy (konradzuse):

wait the chain rule is in there @ agent5x?